Exercise 10-14 Product pricing using total costs LO P1

Steeze Co. makes snowboards and uses the total cost approach in setting product prices. Its costs for producing 16,500 units follow. The company targets a profit of $511,500 on this product.

Variable Costs per Unit Fixed Costs
Direct materials $ 113 Overhead $ 483,000
Direct labor 38 Selling 103,000
Overhead 33 Administrative 338,000
Selling 8


1.
Compute the total cost per unit.
2. Compute the markup percentage on total cost. (Round your final percentage answer to 1 decimal place.)
3. Compute the product’s selling price using the total cost method.

1. Total cost per unit
2. Markup percentage
3. Selling price
Answer #1

1.

Variable Costs per Unit Fixed Costs
Direct materials $ 113 Overhead $ 483,000
Direct labor 38 Selling 103,000
Overhead 33 Administrative 338,000
Selling 8
Total $ 192 Total $ 924,000

Variable Costs per Unit = $192

Total fixed costs = $924,000

Fixed cost per unit = 924,000/16,500

= $56

Total cost per unit = Variable Costs per Unit + Fixed cost per unit

= 192 + 56

= $248

2.

The company targets a profit of $511,500

Profit per unit = Target profit/Number of units

= 511,500/16,500

= $31

Markup percentage = Profit per unit/Total cost per unit

= 31/248

= 12.5%

3.

Selling price per unit = Total cost per unit + Profit per unit

= 248 + 31

= $279

Kindly comment if you need further assistance. Thanks